OpenStudy (anonymous):

For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 2x + y = 8 -x - 3y = -12 3y + 12 -3y + 12 3y - 12 -3y - 12 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

solve -x - 3y = -12 for x

OpenStudy (anonymous):

Where would I multiply that into

jimthompson5910 (jim_thompson5910):

you just need to isolate x (ie get it on one side)

OpenStudy (anonymous):

Divide -12 by 3y?

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

Or er.. by -3y?

jimthompson5910 (jim_thompson5910):

you need to move over the -3y, but you don't divide

OpenStudy (anonymous):

I don't know how I'd do that >. <

jimthompson5910 (jim_thompson5910):

what undoes -3y

OpenStudy (anonymous):

Add?

jimthompson5910 (jim_thompson5910):

yep, add 3y to both sides

OpenStudy (anonymous):

-x = -15 ?

OpenStudy (anonymous):

I think I did that wrong.

jimthompson5910 (jim_thompson5910):

no you don't add 3y to -12 to get -15y

jimthompson5910 (jim_thompson5910):

you just leave it as -x = 3y - 12 so, x = ??

OpenStudy (anonymous):

I still don't understand.

jimthompson5910 (jim_thompson5910):

-x = 3y - 12 x = -3y + 12 ...multiply everything by -1

jimthompson5910 (jim_thompson5910):

so the answer is -3y + 12

OpenStudy (anonymous):

Why did you multiply by -1? Just curious so I can understand what you did.

jimthompson5910 (jim_thompson5910):

well we had -x = 3y - 12 but we want x = ....

jimthompson5910 (jim_thompson5910):

to go from -x to x, you multiply that -x by -1 to get x

jimthompson5910 (jim_thompson5910):

you have to multiply everything by -1 to balance things out

OpenStudy (anonymous):

Ohh, okay.